# Hemlet: A Heterogeneous Compute-in-Memory Chiplet Architecture for Vision Transformers with Group-Level Parallelism

### v2

@@ 4.1. Group-Level Parallelism: Motivation
While static VMMs dominate Transformer operations and consequently consume substantial hardware resources in chiplet systems, conventional layer-wise mapping methods—where weights from a single layer are sequentially mapped to the columns of an ACIM array—often do not fully exploit the available ACIM resources during runtime. As described in Section 2.1, although VMMs are performed in the analog domain, ADCs are necessary to convert the results back for further processing. However, in practical implementations, a single ADC is typically shared across multiple memory columns due to physical layout constraints—specifically, ADCs are significantly wider than the pitch of the RRAM columns (shimengreview). To match the column pitch and reduce ADC area overhead, memory columns are grouped and time-multiplexed via a multiplexer (MUX), allowing only one column per group to access the ADC in each cycle (Fig. 5 (b)). While computation occurs concurrently across all columns, the shared ADC digitization process is inherently serialized through multiplexing, significantly constraining the throughput per subarray, shown in Fig. 5 (c).
@@ 4.2. Cross-layer Interleaved Distribution
Figure 6. Illustration of the interleaved weight distribution strategy. The laye-wise mapping method (left) statically assigns each group to a specific layer, while the interleaved approach (right) distributes weights from multiple layers across groups to enable concurrent group activation.
@@ 4.3. Deployment Approach
    Stage 3: Processing the MHA layers. After completing the slack filling in Stage 2, each type of MHA layer in $\{\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V},\mathbf{W}_{O}\}$ has $R$ remaining layers. We compute $(p,q)=\text{divmod}(R,M)$. If $p\geq 1$, we create $p$ GLP_LayerSets for each of the four MHA types. Moreover, if $q>0$, we handle the $q$ remaining layers of each MHA type as follows: If $3M=4N$: create three GLP_LayerSets for $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$, and evenly distribute the $\mathbf{W}_{O}$ layers among them. Otherwise, all the rest goes to stage 4
@@ 5. System Dataflow Optimization
For the dynamic VMMs operation, the input matrices $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ are partitioned into smaller blocks along the sequence length for blocked processing considering the DCIM capacity limit, adopting the same block size $B_{L}$ as used in the previous stage to match the processing granularity. For the in-memory VMMs, blocks of $\mathbf{Q}$ and $\mathbf{V}$ will be written into the DCIM array as weights for $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ operations separately. To reduce the forward and backward transfers of the intermediate matrix $\mathbf{P}^{\prime}$ between the DCIM chiplet and the IDP, we adopt a placement strategy that prioritizes mapping the PEs responsible for both $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ of the same attention head to the same DCIM chiplet. While the process is blocked, fully executing the Softmax operation between $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ on local DCIM chiplets requires a much larger chiplet buffer to store the entire intermediate matrix $\mathbf{P}^{\prime}$ of size $B_{L}\times L$. To reduce this buffer overhead, we adopt an optimization inspired by FlashAttention (flashattention), a local $\mathrm{Softmax}_{\text{local}}$ is applied to each block result of $\mathbf{P^{\prime}}=\mathbf{Q}\mathbf{K}^{T}$, yielding $\mathbf{P^{\prime\prime}}=\mathrm{Softmax}_{\text{local}}(\mathbf{P^{\prime}})$. The corresponding block output is then obtained as 2 $\mathbf{S^{\prime}}=\mathbf{P^{\prime\prime}}\mathbf{V}$ in Fig. 8 (c). After collecting all $\mathbf{S^{\prime}}$ blocks along the standard Softmax dimension, a 3 global normalization is performed to produce the $\mathbf{S}$ block. In this way, each $\mathbf{P}^{\prime}$ block can be consumed immediately, avoiding much larger on-chip buffer of the DCIM chiplet or the forward and backward transfers between DCIM chiplets and IDP chiplet.
@@ 6.1. Experimental Methodology
Simulator and Modeling Framework: We develop a custom cycle-accurate simulator tailored for heterogeneous CIM-based chiplet systems. The simulation flow integrates hardware-level modeling, dataflow-level scheduling, and event-driven instruction execution across computations on chiplets and inter-chiplet communication. For ACIM chiplets, we adopt NeuroSim (NeuroSimv1.3) to model RRAM-based subarrays and other digital blocks at a 22 nm technology node. DCIM chiplets and the IDP are also extended from NeuroSim and assumed at the 7 nm technology node, with DCIM array performance from the AutoDCIM design (autodcim) and scaled to 7 nm using DeepScale (sarangi2021deepscaletool). To estimate inter-chiplet communication cost, we extend BookSim (jiang2013detailed) for latency estimation under a mesh topology. The corresponding energy consumption is evaluated using post-silicon data from Simba (shao2019simba). To determine the hardware configuration of ACIM, we utilize MICSim (micsim) to evaluate ImageNet-1K(imagenet) classification performance under the ViT-B/16 model. We perform quantization-aware training (QAT) with I-ViT(li2023ivit), achieving $80\%$ accuracy. Our results indicate that with an ADC precision of 9 bits and 2-bit RRAM with an $R{on}/R_{off}$ ratio of 150(rram150), no accuracy degradation is observed compared to our software baseline.
System Configuration: The system is configured to ensure all pretrained weights are fully mapped into ACIM chiplets before inference, consistent with prior works (retransformer; x-former; shafiee2016isaac; chi2016prime; siam). An ACIM chiplet consists of multiple ACIM PEs. Each ACIM PE features a fixed configuration of $60$ ACIM subarrays ($128\times 128$). The group size for the ACIM subarray is set to 8. Meanwhile, the DCIM chiplet comprises multiple DCIM PEs, with each DCIM PE integrating 4 DCIM subarrays ($64\times 64$). To account for the varying capacities of chiplets, we assume that the number of ACIM PEs for the ACIM chiplet and the number of DCIM PEs for the DCIM chiplet may vary. To analyze the impact of chiplet size and communication bandwidth, we evaluate configurations of A18D9 (each ACIM chiplet with 18 ACIM PEs and each DCIM chiplet with 9 DCIM PEs), A32D16, and A50D25 under a range of NoP bandwidths from 8 GB/s to 32 GB/s. The NoP communication network is configured in a 2D mesh topology, which follows the design in Simba(shao2019simba).
@@
Workloads: We evaluate our system using standard ViT models (vit), covering three representative scales: ViT-S/16, ViT-B/16, and ViT-L/16. These models span a wide range of parameter counts and computational requirements, allowing us to assess the scalability and efficiency of our proposed architecture and GLP strategy. All models are evaluated under the image resolution of $224\times 224$, using $16\times 16$ patching, and are fully quantized to 8-bit with I-ViT (li2023ivit) on the ImageNet-1K(imagenet). The detailed model parameters, including embedding dimension, number of Transformer blocks, and the Sequence length, are summarized in Table 1.
@@ 6.3. System-level Performance Analysis
Fig. 10(a) shows that GLP reduces latency in most settings by improving ACIM throughput, with the maximum system speedup reaching $2.53\times$. The effectiveness of this improvement is senstive on both NoP bandwidth and chiplet size. Under lower bandwidth NoP configurations, inter-chiplet communication occupies a larger fraction of the overall execution time, which limits the achievable benefit from ACIM-side acceleration. As the NoP bandwidth increases, the communication bottleneck is gradually alleviated, allowing the throughput advantages of GLP-enabled ACIM execution to translate more effectively into system speedup. The system-level dataflow optimization further improves performance by alleviating inter-chiplet communication, providing greater benefits especially under conditions of limited bandwidth and smaller chiplet sizes. Overall, the proposed Hemlet architecture, which integrates GLP and system-level dataflow optimization, ensures consistent acceleration across all tested configurations, delivering system-level speedups ranging from $1.89\times$ to $4.47\times$.
@@
As shown in Fig.10 (b), we can observe that GLP mapping introduces extra energy overhead in all settings. This arises because, under GLP, the weights of each layer are interleaved across more groups, leaving each ACIM subarray with fewer weight columns. As a result, each input is reused less within one subarray and must be loaded into more subarrays, thereby increasing data movement and overall energy consumption. After applying the system-level dataflow optimization, the overall energy consumption can be reduced. This is because the optimization decreases inter-chiplet communication, thereby lowering the energy cost of data transfers. As a result, the system-level dataflow optimization can partially compensate for the additional energy overhead introduced by GLP. Still, the remaining energy overhead is around $10\%$.
@@
Comparison with Existing Work: In Table 2, we compare our design with two state-of-the-art ViT accelerators. The reported results for Hemlet correspond to the configuration using the A32D16 chiplet setup with 32 GB/s NoP bandwidth, evaluated on the ViT-L/16 model for end-to-end inference, where each processed element in non-VMM operations is counted as one operation. The work presented in (dac24dcim) utilizes DCIM arrays as the computation cores for both static VMMs and dynamic VMMs. While the reload capability of the DCIM computation units allows for a monolithic design with limited on-chip resources, this approach results in relatively low throughput due to the necessary weight reloads. On the other hand, H3DAtten (li2023h3datten) shares the same design principles as Hemlet, utilizing heterogeneous computing units and employing ACIM arrays to store all weights on-chip. To address area constraints, they employ 3D integration of the eNVM arrays with the other digital circuits. This 3D integration mitigates technology lock-in, as the top and bottom tiers can be fabricated using different process nodes. Hemlet achieves superior throughput compared to H3DAtten, which can be mainly attributed to two key reasons. First, H3DAtten employs a lower precision ADC, necessitating a reduction in row parallelism in the CIM computation to maintain accuracy. Second, Hemlet enhances throughput further by implementing GLP and optimizing dataflow at the system-level. H3DAtten demonstrates better energy efficiency than Hemlet, which can also be attributed to two primary factors. First, the 3D integration in H3DAtten allows for a simpler NoC and incurs lower communication overhead compared to the NoP adopted by Hemlet. Second, H3DAtten locates its ADCs on the bottom tier and fabricates them using a more advanced technology node. In contrast, Hemlet integrates the ADCs on the ACIM chiplet, sharing the same technology node as the eNVM arrays. While this 3D integration in H3DAtten enhances energy efficiency, it also introduces greater design complexity and reduces flexibility due to the tight coupling between the top-tier array and the bottom-tier ADC.
@@
Work DAC’24(dac24dcim) H3DAtten(li2023h3datten) Hemlet
@@ 7. Conclusion
Acknowledgements.
This work is supported in part by Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0540 ), Guangzhou Municipal Science and Technology Project (Municipal Key Laboratory Construction Project, Grant No.2023A03J0013), and ”1+1+1” Joint Funding Program (No. JC031).

### v3

@@ 4.1. Group-Level Parallelism: Motivation
While static VMMs dominate Transformer operations and consequently consume substantial hardware resources in chiplet systems, conventional layer-wise mapping methods—where weights from a single layer are sequentially mapped to the columns of an ACIM array—often do not fully exploit the available ACIM resources during runtime. As described in Section 2.1, although VMMs are performed in the analog domain, ADCs are necessary to convert the results back for further processing. However, in practical implementations, a single ADC is typically shared across multiple memory columns due to physical layout constraints—specifically, ADCs are significantly wider than the pitch of the RRAM columns (Yu et al., 2021b). To match the column pitch and reduce ADC area overhead, memory columns are grouped and time-multiplexed via a multiplexer (MUX), allowing only one column per group to access the ADC in each cycle (Fig. 4 (b)). While computation occurs concurrently across all columns, the shared ADC digitization process is inherently serialized through multiplexing, significantly constraining the throughput per subarray, shown in Fig. 4 (c).
@@ 4.2. Cross-layer Interleaved Distribution
Figure 5. Illustration of the interleaved weight distribution strategy. The layer-wise mapping method (left) statically assigns each group to a specific layer, while the interleaved approach (right) distributes weights from multiple layers across groups to enable concurrent group activation.
@@ 4.3. Deployment Approach
    Stage 3: Processing the MHA layers. After completing the slack filling in Stage 2, each type of MHA layer in $\{\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V},\mathbf{W}_{O}\}$ has $R$ remaining layers. We compute $(p,q)=\text{divmod}(R,M)$. If $p\geq 1$, we create $p$ GLP_LayerSets for each of the four MHA types. Moreover, if $q>0$, we handle the $q$ remaining layers of each MHA type as follows: If $3M=4N$: create three GLP_LayerSets for $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$, and evenly distribute the $\mathbf{W}_{O}$ layers among them. Otherwise, all the rest goes to stage 4.
@@ 5. System Dataflow Optimization
For the dynamic VMMs operation, the input matrices $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ are partitioned into smaller blocks along the sequence length for blocked processing considering the DCIM capacity limit, adopting the same block size $B_{L}$ as used in the previous stage to match the processing granularity. For the in-memory VMMs, blocks of $\mathbf{Q}$ and $\mathbf{V}$ will be written into the DCIM array as weights for $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ operations separately. To reduce the forward and backward transfers of the intermediate matrix $\mathbf{P}^{\prime}$ between the DCIM chiplet and the IDP, we adopt a placement strategy that prioritizes mapping the PEs responsible for both $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ of the same attention head to the same DCIM chiplet. While the process is blocked, fully executing the Softmax operation between $\mathbf{Q}\mathbf{K}^{T}$ and $\mathbf{P}\mathbf{V}$ on local DCIM chiplets requires a much larger chiplet buffer to store the entire intermediate matrix $\mathbf{P}^{\prime}$ of size $B_{L}\times L$. To reduce this buffer overhead, we adopt an optimization inspired by FlashAttention (Dao et al., 2022), a local $\mathrm{Softmax}_{\text{local}}$ is applied to each block result of $\mathbf{P^{\prime}}=\mathbf{Q}\mathbf{K}^{T}$, yielding $\mathbf{P^{\prime\prime}}=\mathrm{Softmax}_{\text{local}}(\mathbf{P^{\prime}})$. The corresponding block output is then obtained as 2 $\mathbf{S^{\prime}}=\mathbf{P^{\prime\prime}}\mathbf{V}$ in Fig. 7 (c). After collecting all $\mathbf{S^{\prime}}$ blocks along the standard Softmax dimension, a 3 global normalization is performed to produce the $\mathbf{S}$ block. In this way, each $\mathbf{P}^{\prime}$ block can be consumed immediately, avoiding much larger on-chip buffer of the DCIM chiplet or the forward and backward transfers between DCIM chiplets and IDP chiplet.
@@ 6.1. Experimental Methodology
Simulator and Modeling Framework: We develop a custom cycle-accurate simulator tailored for heterogeneous CIM-based chiplet systems. The simulation flow integrates hardware-level modeling, dataflow-level scheduling, and event-driven instruction execution across computations on chiplets and inter-chiplet communication. For ACIM chiplets, we adopt NeuroSim (Peng et al., 2019) to model RRAM-based subarrays and other digital blocks at a 22 nm technology node. DCIM chiplets and the IDP are also extended from NeuroSim and assumed at the 7 nm technology node, with DCIM array performance from the AutoDCIM design (Chen et al., 2023) and scaled to 7 nm using DeepScale (Sarangi and Baas, 2021). To estimate inter-chiplet communication cost, we extend BookSim (Jiang et al., 2013) for latency estimation under a mesh topology. The corresponding energy consumption is estimated based on the advanced packaging parameters provided in UCIe (Sharma, 2022). To determine the hardware configuration of ACIM, we utilize MICSim (Wang et al., 2025a) to evaluate ImageNet-1K(Deng et al., 2009) classification performance under the ViT-B/16 model. We perform quantization-aware training (QAT) with I-ViT(Li and Gu, 2023), achieving $80\%$ accuracy. Our results indicate that with an ADC precision of 9 bits and 2-bit RRAM with an $R_{on}/R_{off}$ ratio of 150(Yin et al., 2019), no accuracy degradation is observed compared to our software baseline.
System Configuration: The system is configured to ensure all pretrained weights are fully mapped into ACIM chiplets before inference, consistent with prior works (Yang et al., 2020; Sridharan et al., 2023; Shafiee et al., 2016; Chi et al., 2016; Krishnan et al., 2021). An ACIM chiplet consists of multiple ACIM PEs. Each ACIM PE features a fixed configuration of $60$ ACIM subarrays ($128\times 128$). The group size for the ACIM subarray is set to 8. The buffer size of each ACIM chiplet is 64 KB. Meanwhile, the DCIM chiplet comprises multiple DCIM PEs, with each DCIM PE integrating 4 DCIM subarrays ($64\times 64$). The buffer size of each DCIM chiplet is 512 KB. To account for the varying capacities of chiplets, we assume that the number of ACIM PEs for the ACIM chiplet and the number of DCIM PEs for the DCIM chiplet may vary. To analyze the impact of chiplet size and communication bandwidth, we evaluate configurations of A18D9 (each ACIM chiplet with 18 ACIM PEs and each DCIM chiplet with 9 DCIM PEs), A32D16, and A50D25 under a range of NoP bandwidths from 8 GB/s to 32 GB/s. The NoP communication network is configured in a 2D mesh topology, which follows the design in Simba(Shao et al., 2019).
@@
Workloads: We evaluate our system using standard ViT models (Dosovitskiy et al., 2021), covering three representative scales: ViT-S/16, ViT-B/16, and ViT-L/16. These models span a wide range of parameter counts and computational requirements, allowing us to assess the scalability and efficiency of our proposed architecture and GLP strategy. All models are evaluated under the image resolution of $224\times 224$, using $16\times 16$ patching, and are fully quantized to 8-bit with I-ViT (Li and Gu, 2023) on the ImageNet-1K(Deng et al., 2009). The detailed model parameters, including embedding dimension, number of Transformer blocks, and the Sequence length, are summarized in Table 1.
@@ 6.3. System-level Performance Analysis
Fig. 9(a) shows that GLP reduces latency in most settings by improving ACIM throughput, with the maximum system speedup reaching $2.53\times$. The effectiveness of this improvement is sensitive on both NoP bandwidth and chiplet size. Under lower bandwidth NoP configurations, inter-chiplet communication occupies a larger fraction of the overall execution time, which limits the achievable benefit from ACIM-side acceleration. As the NoP bandwidth increases, the communication bottleneck is gradually alleviated, allowing the throughput advantages of GLP-enabled ACIM execution to translate more effectively into system speedup. The system-level dataflow optimization further improves performance by alleviating inter-chiplet communication, providing greater benefits especially under conditions of limited bandwidth and smaller chiplet sizes. Overall, the proposed Hemlet architecture, which integrates GLP and system-level dataflow optimization, ensures consistent acceleration across all tested configurations, delivering system-level speedups ranging from $2.41\times$ to $5.74\times$.
@@
As shown in Fig.9 (b), we can observe that GLP mapping introduces extra energy overhead in all settings. This arises because, under GLP, the weights of each layer are interleaved across more groups, leaving each ACIM subarray with fewer weight columns. As a result, each input is reused less within one subarray and must be loaded into more subarrays, thereby increasing data movement and overall energy consumption. After applying the system-level dataflow optimization, the overall energy consumption can be reduced. This is because the optimization decreases inter-chiplet communication, thereby lowering the energy cost of data transfers. As a result, the system-level dataflow optimization can partially compensate for the additional energy overhead introduced by GLP, with an average overhead of $4.62\%$.
@@
Comparison with Existing Work: In Table 2, we compare our design with two state-of-the-art ViT accelerators. The reported results for Hemlet correspond to the configuration using the A32D16 chiplet setup with 32 GB/s NoP bandwidth, evaluated on the ViT-L/16 model for end-to-end inference, where each processed element in non-VMM operations is counted as one operation. The work presented in (Chen et al., 2024) utilizes DCIM arrays as the computation cores for both static VMMs and dynamic VMMs. While the reload capability of the DCIM computation units allows for a monolithic design with limited on-chip resources, this approach results in relatively low throughput due to the necessary weight reloads. On the other hand, H3DAtten (Li et al., 2023) shares the same design principles as Hemlet, utilizing heterogeneous computing units and employing ACIM arrays to store all weights on-chip. To address area constraints, they employ 3D integration of the eNVM arrays with the other digital circuits. This 3D integration mitigates technology lock-in, as the top and bottom tiers can be fabricated using different process nodes. Hemlet achieves superior throughput compared to H3DAtten, which can be mainly attributed to two key reasons. First, H3DAtten employs a lower precision ADC, necessitating a reduction in row parallelism in the CIM computation to maintain accuracy. Second, Hemlet enhances throughput further by implementing GLP and optimizing dataflow at the system-level. H3DAtten demonstrates better energy efficiency than Hemlet, which can also be attributed to two primary factors. First, the 3D integration in H3DAtten allows for a simpler NoC and incurs lower communication overhead compared to the NoP adopted by Hemlet. Second, H3DAtten locates its ADCs on the bottom tier and fabricates them using a more advanced technology node. In contrast, Hemlet integrates the ADCs on the ACIM chiplet, sharing the same technology node as the eNVM arrays. While this 3D integration in H3DAtten enhances energy efficiency, it also introduces greater design complexity and reduces flexibility due to the tight coupling between the top-tier array and the bottom-tier ADC.
@@
Work DAC’24(Chen et al., 2024) H3DAtten(Li et al., 2023) Hemlet